# HIGHT: Hierarchical Graph Tokenization for Graph-Language Alignment

### v1

@@ Abstract
Recently there has been a surge of interest in extending the success of large language models (LLMs) to graph modality, such as social networks and molecules. As LLMs are predominantly trained with 1D text data, most existing approaches adopt a graph neural network to represent a graph as a series of node tokens and feed these tokens to LLMs for graph-language alignment. Despite achieving some successes, existing approaches have overlooked the hierarchical structures that are inherent in graph data. Especially, in molecular graphs, the high-order structural information contains rich semantics of molecular functional groups, which encode crucial biochemical functionalities of the molecules. We establish a simple benchmark showing that neglecting the hierarchical information in graph tokenization will lead to subpar graph-language alignment and severe hallucination in generated outputs. To address this problem, we propose a novel strategy called HIerarchical GrapH Tokenization (HIGHT). HIGHT employs a hierarchical graph tokenizer that extracts and encodes the hierarchy of node, motif, and graph levels of informative tokens to improve the graph perception of LLMs. HIGHT also adopts an augmented graph-language supervised fine-tuning dataset, enriched with the hierarchical graph information, to further enhance the graph-language alignment. Extensive experiments on $7$ molecule-centric benchmarks confirm the effectiveness of HIGHT in reducing hallucination by $40$%, as well as significant improvements in various molecule-language downstream tasks.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in understanding and processing natural languages (Radford et al., 2019; OpenAI, 2022; Touvron et al., 2023a; Bubeck et al., 2023). Recently, there has been a surge of interest in extending the capabilities of LLMs to graph modality (Jin et al., 2023; Li et al., 2023d; Wei et al., 2024; Mao et al., 2024; Fan et al., 2024) such as social networks (Tang et al., 2023; Chen et al., 2024) and molecular graphs (Liu et al., 2023d; Zhao et al., 2023; Cao et al., 2023; Li et al., 2024). Inspired by the success of large vision-language models (Zhang et al., 2024; Zhu et al., 2023; Liu et al., 2023b), existing large graph-language models (LGLMs) predominantly adopt a graph neural network (GNN) (Kipf and Welling, 2017; Hamilton et al., 2017; Xu et al., 2019) to tokenize graph information as a series of node embeddings (or node tokens), and then leverage an adapter such as a Multi-layer perceptron (MLP) or a Q-former (Li et al., 2023a) to transform the node tokens into those compatible with LLMs (Fan et al., 2024). To facilitate the alignment of graph and language modalities, LGLMs will undergo a graph-language instruction tuning stage with the graph and the corresponding caption data, so as to realize the graph-language alignment (Jin et al., 2023; Li et al., 2023d; Fan et al., 2024).
@@
Figure 1: Illustration of HIGHT. (a) Given a molecule (i.e., PubChem ID 3, 5,6-Dihydroxycyclohexa-1,3-diene-1-carboxylic acid), HIGHT detects the motifs and incorporates the “supernodes” for each motif (The whole graph is also considered as a “super motif”.). Then, HIGHT tokenizes the molecule into both node-level (i.e., atoms) and motif-level (i.e., functional groups) tokens. The hierarchical view enables LLMs to align the molecular structures and the language descriptions of the molecule better. (b) Therefore, HIGHT significantly reduces the hallucination of LGLMs and improves the downstream performance across various molecule-centric tasks. All metrics are transformed a bit such that a higher number means a better downstream task performance.
Despite achieving certain successes, the graph tokenization in existing LGLMs neglects the essential hierarchical structures that are inherent in graph data (Ying et al., 2018). Especially, in molecular graphs, the high-order structural information, such as motifs or functional groups, contains rich semantics of the biochemical functionalities of the molecules (Milo et al., 2002; Bohacek et al., 1996; Sterling and Irwin, 2015). For example, the existence of the hydroxide functional group (“-OH”) in small molecules often indicates a higher water solubility. Therefore, the perception of functional groups in a molecule is essential for LLMs to perceive the molecules. Intuitively, feeding LLMs with only node tokens makes the molecule understanding harder as LLMs have to learn to combine atoms in a functional group during the instruction tuning phase. However, atoms are usually treated as separate tokens in LGLMs and there is often a lack of supervision signal to prompt about the combinations of specific motifs. Consequently, neglecting the hierarchical information will lead to subpar graph-language alignment and severe hallucination. To demonstrate the issue, we construct a simple benchmark called MotifHallu that queries LLMs about the existence of common functional groups. Surprisingly, we find that existing LGLMs consistently answer “Yes” when asking about any functional groups, as demonstrated in Sec. 3.2. It then raises a challenging research question:
@@
Is there a way to incorporate the intrinsic hierarchical graph information into LLMs?
@@
In this paper, we study the problem with a focus on molecular data, and introduce a new graph-language alignment strategy called HIerarchical GrapH Tokenization (HIGHT). As shown in Fig. 1, HIGHT includes a hierarchical graph tokenizer, as well as a hierarchical molecular instruction tuning dataset to facilitate a better alignment of molecule and language modalities. Inspired by the success of hierarchical GNNs in molecular representation learning (ZHANG et al., 2021; Zang et al., 2023; Inae et al., 2023; Luong and Singh, 2023), we transform the original molecular graph into a hierarchical graph with motif and graph nodes added in. Then, we employ a Vector Quantized-Variational AutoEncoder (VQVAE) to obtain atom-level, motif-level, and graph-level tokens separately with the self-supervised tasks (Zang et al., 2023). To retain more original structural information, we further attach Laplacian positional encodings to the tokens. After that, we adopt a multi-level adapter consisting of three adapters processing atom-level, motif-level, and graph-level tokens, respectively, before feeding them into the LLMs. In addition, to facilitate the use of hierarchical information encoded by the tokens, we augment the original molecular instruction tuning dataset with motif descriptions. Our contributions can be summarized as follows:
@@
  • To the best of our knowledge, we are the first to propose incorporating the hierarchical graph information into LGLMs with new architectures and instruction tuning dataset HiPubChem.
@@
  • To facilitate the graph-language alignment study on molecular graphs, we further propose the first hallucination benchmark MotifHallu based on the existence of common functional groups.
@@
  • We conduct extensive experiments with $7$ real-world molecular and reaction comprehension benchmarks. The results show that HIGHT significantly reduces the hallucination issue on MotifHallu by up to $40$% and consistently improves the downstream molecule-language performances.
@@ 2 Preliminaries
We begin by introducing preliminary concepts and related works of LGLMs.
Large Graph-Language Models. As LLMs have demonstrated great capabilities across a wide range of natural language tasks, there has been an increasing interest in extending LLMs to broader applications where the text data are associated with the structure information (i.e., graphs) (Jin et al., 2023; Li et al., 2023d; Wei et al., 2024; Mao et al., 2024; Fan et al., 2024). A graph can be denoted as ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ with a set of $n$ nodes $v\in{\mathcal{V}}$ and a set of $m$ edges $(u,v)\in{\mathcal{E}}$. Each node $u$ has node attributes as ${\bm{x}}_{u}\in\mathbb{R}^{d}$ and each edge $(u,v)$ has edge attributes $e_{u,v}\in\mathbb{R}^{d_{e}}$. A number of LGLMs have been developed to process graph-text associated dataset ${\mathcal{D}}=\{{\mathcal{G}},{\bm{c}}\}$, where ${\bm{c}}=[c_{1},...,c_{l_{c}}]$ refers to the caption of the graph ${\mathcal{G}}$. For node-centric tasks, ${\bm{c}}_{i}$ will associate with the nodes (Tang et al., 2023), while in this paper we focus on graph-centric tasks, i.e., molecules and molecular captions (Liu et al., 2023d). Usually, an $l$-layer GNN is employed to encode a graph as:
${\bm{h}}^{(l)}_{u}=\text{COMBINE}({\bm{h}}^{(l-1)}_{u},\text{AGG}(\{({\bm{h}}^{(l-1)}_{u},{\bm{h}}^{(l-1)}_{v},e_{uv})|v\in{\mathcal{N}}(u)\})),$ (1)
where ${\bm{h}}^{(l)}_{u}\in\mathbb{R}^{h}$ refers to the node embedding of node $u$ after $l$ layers of GNN, $\text{AGG}(\cdot)$ is the aggregation function (e.g., mean) among the information from neighbors of node $u$, and COMBINE is the operator for combining information of node $u$ with its neighbors ${\mathcal{N}}(u)$ (e.g., concatenation). Then, after $l$ message passing iterations, the graph-level embedding can be obtained as:
@@
where $\text{READOUT}(\cdot)$ is a pooling operator (e.g., mean pooling) among all the node embeddings. With the representations of the nodes and graphs, LGLMs can fuse the graph and language information in various ways, such as transforming into natural languages describing the graphs (Fatemi et al., 2024), or neural prompts within the LLMs (Tian et al., 2024). In addition, the embeddings can also be leveraged to postprocess the LLM outputs (Liu et al., 2024a). Orthogonal to different fusion mechanisms, in this work, we will focus on transforming graph embeddings into input tokens of LLMs to demonstrate the benefits of hierarchical graph modeling, which can be formulated as (Tang et al., 2023; Chen et al., 2024; Liu et al., 2023d; Zhao et al., 2023; Cao et al., 2023; Li et al., 2024):
$p_{\mathbf{\theta}}({\bm{a}}|{\bm{q}},{\bm{h}})=\Pi_{i=1}^{l_{a}}p_{\mathbf{\theta}}({\bm{a}}_{i}|{\bm{q}},f_{n}({\bm{h}}),{\bm{a}}_{<i}),$ (3)
@@
where the LGLM is required to approximate $p_{\mathbf{\theta}}$ to output the desired answer ${\bm{a}}$ given the question ${\bm{q}}$, and the graph tokens ${\bm{h}}$ adapted with adapter $f_{n}:\mathbb{R}^{h}\rightarrow\mathbb{R}^{h_{e}}$ that projects the graph tokens to the embedding space of LLMs. In addition, one could also incorporate the 1D sequence of molecules such as SMILES (Weininger, 1988) and SELFIES (Krenn et al., 2019) into ${\bm{q}}$ and ${\bm{a}}$ to facilitate the alignment.
@@
Molecular Foundation Models. More specifically, this work focuses on one of the most popular graph-language alignment tasks, i.e., molecule-language alignment (Liu et al., 2024c; Pei et al., 2024). In fact, there is a separate line of works aiming to develop language models for molecules and proteins – the language of lives, from 1D sequences such as SMILES (Irwin et al., 2022), 2D molecular graphs (Wang et al., 2022), 3D geometric conformations (Liu et al., 2022; Zhou et al., 2023), to scientific text (Beltagy et al., 2019) and multimodal molecule-text data (Liu et al., 2023c; Luo et al., 2023a; Christofidellis et al., 2023; Liu et al., 2024b; Su et al., 2022; Zeng et al., 2022). The adopted backbones range from encoder-decoder architectures such as MolT5 (Edwards et al., 2022) and Galactica (Taylor et al., 2022), to auto-regressive language modeling (Luo et al., 2023b; Liu et al., 2023f). Inspired by the success of large vision-language models (Li et al., 2023a; Zhu et al., 2023; Liu et al., 2023b), the community further seeks for developing molecular foundation models built upon existing molecular language models with more sophisticated graph information fusion modules. For example, Liu et al. (2023d); Zhao et al. (2023) develop advanced cross-modal adapters and generalized position embeddings to promote a better graph-language alignment of encoder-decoder based molecular foundation models. Liang et al. (2023); Cao et al. (2023); Li et al. (2024) develop cross-modal adapters for decoder only language models such as Llama (Touvron et al., 2023a). Orthogonal to the aforementioned works, we focus more on what information one shall extract from the graph for better graph-language alignment. We choose to build our methods upon decoder only language models, with the hope to build a versatile agent that can perceive graph information beyond the language, image, and audio modalities (Xi et al., 2023).
@@
Hierarchical Graph Representation Learning. The hierarchical nature has been widely and explicitly incorporated in learning high-quality graph representations (Ying et al., 2018). Especially in molecular graphs, the high-order structural information naturally captures the existence of motifs and functional groups. Therefore, the hierarchy of node-motif-graph has been widely applied in self-supervised molecular representation learning (ZHANG et al., 2021; Zang et al., 2023; Inae et al., 2023; Luong and Singh, 2023). Nevertheless, it remains unclear how to properly incorporate the hierarchical graph information in graph instruction tuning with LLMs.
3 Graph Tokenization for Graph-Language Alignment
Existing LGLMs predominantly tokenize a graph into a series of node embeddings (or node tokens). Despite achieving some success, we find that node-centric tokenization can hardly express motif-level information and will lead to issues such as hallucination as demonstrated in this section.

### v2

# Hierarchical Graph Tokenization for Molecule-Language Alignment

@@ Abstract
Recently, there has been a surge of interest in extending the success of large language models (LLMs) from texts to molecules. Most existing approaches adopt a graph neural network to represent a molecule as a series of node tokens for molecule-language alignment, which, however, have overlooked the inherent hierarchical structures in molecules. Notably, higher-order molecular structures contain rich semantics of functional groups, which encode crucial biochemical functionalities of the molecules. We show that neglecting the hierarchical information in tokenization will lead to subpar molecule-language alignment and severe hallucination. To address this limitation, we propose HIerarchical GrapH Tokenization (HIGHT). HIGHT employs a hierarchical graph tokenizer that encodes the hierarchy of atom, motif, and molecular levels of informative tokens to improve the molecular perception of LLMs. HIGHT also adopts an augmented instruction tuning dataset, enriched with the hierarchical graph information, to further enhance the molecule-language alignment. Extensive experiments on $14$ real-world benchmarks verify the effectiveness of HIGHT in reducing hallucination by $40$%, and significant improvements in various molecule-language downstream tasks. The project is available at https://higraphllm.github.io/.
Large Language Models, Graph Neural Networks, Molecule Understanding, Alignment, Tokenization
[…]
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in understanding and processing natural languages (Radford et al., 2019; OpenAI, 2022; Touvron et al., 2023a; Bubeck et al., 2023). Recently, there has been a surge of interest in extending the capabilities of LLMs to graph-structured data (Jin et al., 2023; Li et al., 2023d; Wei et al., 2024; Mao et al., 2024; Fan et al., 2024), particularly molecular graphs (Zhao et al., 2023; Cao et al., 2023). Inspired by the success of large vision-language models (Zhang et al., 2024b; Liu et al., 2023a), recent efforts in developing large graph-language models (LGLMs) typically adopt a graph neural network (GNN) (Xu et al., 2019) to tokenize molecules as a series of node embeddings (or node tokens), and then leverage an adapter such as a Multi-layer perceptron (MLP) or a Q-former (Li et al., 2023a) to transform the node tokens into those compatible with LLMs (Fan et al., 2024). To bridge the gap between the graph and language modalities, LGLMs will undergo a molecule-language instruction tuning with the molecular graph and the corresponding captions describing the molecules (Jin et al., 2023; Li et al., 2023d; Fan et al., 2024).
@@
Figure 1: (a) Illustration of HIGHT: Given a molecule (i.e., PubChem ID 3, 5,6-Dihydroxycyclohexa-1,3-diene-1-carboxylic acid), HIGHT detects the motifs and incorporates the “supernodes” for each motif (The whole graph is also considered as a “super motif”.). Then, HIGHT tokenizes the molecule into both node-level (i.e., atoms) and motif-level (i.e., functional groups) tokens. The hierarchical view enables LLMs to align the molecular structures and the language descriptions of the molecule better. (b) Performance Overview: HIGHT significantly reduces the hallucination of LGLMs and improves the downstream performance across various molecule-centric tasks. Due to the heterogeneity of the evaluation metrics in each task, we perform some transformations on the numerical values. In MotifHallu, we report the macro F1 scores. For Property Classification and Molecular Caption, we report the averaged scores of all the subtasks or submetrics. For Property Regression, we normalize the values to the range between $1$ and $100$, i.e., for $a$, the reported number is $0.5/a$. For Chemical Reaction Prediction, we report the averaged values of BLEU, RDK, MACCS, and MORGAN.
Despite recent progress, the tokenization in existing LGLMs neglects the essential hierarchical structures inherent in molecular graphs. In particular, in molecular graphs, the high-order substructures, such as motifs or functional groups, encode rich semantics of the biochemical functionalities of the molecules (Milo et al., 2002; Bohacek et al., 1996; Sterling & Irwin, 2015). For example, the presence of a hydroxide functional group (“-OH”) often indicates a higher water solubility. Therefore, such substructural cues are essential for enabling LLMs to reason about the molecules in a chemically meaningful way. However, existing LGLMs mostly tokenize molecules solely at the atom (node) level, and feed LLMs with only node-level tokens. Consequently, it requires LLMs to implicitly infer the underlying substructures during the instruction tuning stage. The absence of the critical substructures not only increases the unnecessary burdern on the LLMs, but also leads to misaligned representations and a higher likelihood of hallucinations in downstream tasks. To quantify the issue, we introduce a diagnostic benchmark, called MotifHallu, which evaluates the perception ability of LGLMs about the existence of common functional groups. Surprisingly, we find that existing LGLMs often produce false-positive predictions (i.e., keep answering “Yes” for any functional groups), highlighting a critical limitation in current graph tokenization strategies (Sec. 3.2). This observation motivates the following research question:
@@
Is there a feasible approach to integrate the intrinsic hierarchical molecular information into LLMs?
@@
To tackle the problem, we propose a new molecule-language alignment strategy called HIerarchical GrapH Tokenization (HIGHT). As illustrated in Fig. 1, HIGHT adopts a hierarchical graph tokenizer and a hierarchical molecular instruction tuning dataset to facilitate a better alignment of molecule and language modalities. Specifically, inspired by the success of hierarchical GNNs in molecular representation learning (Zhang et al., 2021; Zang et al., 2023; Inae et al., 2023; Luong & Singh, 2023), HIGHT transforms the original molecular graph into a hierarchical graph with motif-level and molecule-level nodes added in. Then, HIGHT employs a Vector Quantized-Variational AutoEncoder (VQVAE) to obtain atom-level, motif-level, and molecule-level tokens separately with the self-supervised tasks (Zang et al., 2023).
In addition, to further encourage the encoding and alignment of hierarchical information, HIGHT augments the original molecular instruction tuning dataset with motif-level descriptions. Our contributions can be summarized as follows:
@@
  • To the best of our knowledge, we are the first to incorporate the hierarchical graph information into LGLMs, with the consideration of both the architecture-level and the instruction tuning data.
@@
  • To facilitate the molecule-language alignment study, we also propose the first hallucination benchmark MotifHallu, synthesized through question-answering based on common functional groups.
@@
  • We conduct extensive experiments with $14$ real-world benchmarks. The results show that HIGHT significantly reduces the hallucination on MotifHallu by up to $40$% and consistently improves the performances on downstream molecule-language tasks.
Hence, HIGHT together with MotifHallu and HiPubChem, lay the solid foundation for developing graph foundation models via graph-language alignment.
@@ 2 Preliminaries
Large Graph-Language Models. As LLMs have demonstrated great capabilities across a wide range of natural language tasks, there has been an increasing interest in extending LLMs to broader applications where the text data are associated with the structure information (i.e., graphs) (Jin et al., 2023; Li et al., 2023d; Wei et al., 2024; Mao et al., 2024; Fan et al., 2024). A graph can be denoted as ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ with a set of $n$ nodes $v\in{\mathcal{V}}$ and a set of $m$ edges $(u,v)\in{\mathcal{E}}$. Each node $u$ has node attributes as ${\bm{x}}_{u}\in\mathbb{R}^{d}$ and each edge $(u,v)$ has edge attributes $e_{u,v}\in\mathbb{R}^{d_{e}}$. A number of LGLMs have been developed to process graph-text associated data ${\mathcal{D}}=\{{\mathcal{G}},{\bm{c}}\}$, where ${\bm{c}}=[c_{1},...,c_{l_{c}}]$ is to the caption of the graph ${\mathcal{G}}$. For node-centric tasks, ${\bm{c}}_{i}$ will associate with the nodes (Tang et al., 2023), while in this paper we focus on graph-centric tasks, i.e., molecules and molecular captions (Liu et al., 2023c). Usually, an $l$-layer GNN is employed to encode a graph as:
${\bm{h}}^{(l)}_{u}=\text{COM}({\bm{h}}^{(l-1)}_{u},\text{AGG}(\{({\bm{h}}^{(l-1)}_{u},{\bm{h}}^{(l-1)}_{v})|v\in{\mathcal{N}}(u)\})),$ (1)
where ${\bm{h}}^{(l)}_{u}\in\mathbb{R}^{h}$ refers to the node embedding of node $u$ after $l$ layers of GNN, $\text{AGG}(\cdot)$ is the aggregation function (e.g., mean) among the information from neighbors of node $u$, and COM is the operator for combining information of node $u$ with its neighbors ${\mathcal{N}}(u)$ (e.g., concatenation). Then, after $l$ message passing iterations, the graph-level embedding can be obtained as:
@@
where $\text{READOUT}(\cdot)$ is a pooling operator (e.g., mean pooling) among all the node embeddings. With the representations of the nodes and graphs, LGLMs can fuse the graph and language information in various ways, such as transforming into natural languages describing the graphs (Fatemi et al., 2024), or neural prompts within the LLMs (Tian et al., 2024). In addition, the embeddings can also be leveraged to post-process the LLM outputs (Liu et al., 2024b). Orthogonal to different fusion mechanisms, in this work, we focus on transforming graph embeddings into input tokens of LLMs, which can be formulated as (Tang et al., 2023; Chen et al., 2024a; Liu et al., 2023c; Zhao et al., 2023; Cao et al., 2023; Li et al., 2024):
$p_{\mathbf{\theta}}({\bm{a}}|{\bm{q}},{\bm{h}})=\text{$\prod$}_{i=1}^{l_{a}}p_{\mathbf{\theta}}({\bm{a}}_{i}|{\bm{q}},f_{n}({\bm{h}}),{\bm{a}}_{<i}),$ (3)
@@
where the LGLM is required to approximate $p_{\mathbf{\theta}}$ to output the desired answer ${\bm{a}}$ given the question ${\bm{q}}$, and the graph tokens ${\bm{h}}$ adapted with adapter $f_{n}:\mathbb{R}^{h}\rightarrow\mathbb{R}^{h_{e}}$ that projects the graph tokens to the embedding space of LLMs. One could also incorporate the 1D information such as SMILES (Weininger, 1988) into ${\bm{q}}$ and ${\bm{a}}$ for alignment.
@@
Molecular Foundation Models. There is a separate line of works aiming to develop language models for molecules and proteins – the language of lives, from 1D sequences such as SMILES (Irwin et al., 2022), 2D molecular graphs (Rong et al., 2020; Wang et al., 2022; Zhang et al., 2024a), 3D geometric conformations (Liu et al., 2022; Zhou et al., 2023), to scientific text (Beltagy et al., 2019) and multimodal molecule-text data (Liu et al., 2023b; Luo et al., 2023a; Christofidellis et al., 2023; Liu et al., 2024c; Su et al., 2022; Zeng et al., 2022; Srinivas & Runkana, 2024). The adopted backbones range from encoder-decoder architectures such as MolT5 (Edwards et al., 2022) and Galactica (Taylor et al., 2022), to auto-regressive language modeling (Luo et al., 2023b; Liu et al., 2023e). Inspired by the success of large vision-language models (Li et al., 2023a; Zhu et al., 2023; Liu et al., 2023a), the community further seeks to develop molecular foundation models built upon existing molecular language models with more sophisticated graph information fusion modules. For example, Liu et al. (2023c); Zhao et al. (2023) develop advanced cross-modal adapters and generalized position embeddings to promote better alignment based on encoder-decoder-based molecular language models. Liang et al. (2023); Cao et al. (2023); Li et al. (2024) develop cross-modal adapters for decoder only language models such as Llama (Touvron et al., 2023a). Orthogonal to the aforementioned works, we focus more on what information one shall extract from the molecules for the alignment. We choose to build our methods upon decoder-only language models, with the hope of building a versatile agent that can perceive molecules beyond the language, image, and audio modalities (Xi et al., 2023).
In the meantime, existing works also try to enrich the molecule-language alignment with additional modalities, such as 2D (Liu et al., 2023c) and 3D (Li et al., 2024) information. In contrast, we focus on the intrinsic hierarchical information of the molecules, such as motifs.
@@
Hierarchical Graph Representation Learning. The hierarchical nature has been widely incorporated in learning high-quality graph representations (Ying et al., 2018). Especially in molecular graphs, the high-order structural information naturally captures the existence of motifs and functional groups. Therefore, the hierarchy of atom-motif-molecule has been widely applied in self-supervised molecular representation learning (Zhang et al., 2021; Zang et al., 2023; Inae et al., 2023; Luong & Singh, 2023). Nevertheless, how to properly incorporate the hierarchical information in molecular instruction tuning of LGLMs remains unclear.
In addition, concurrent works by Park et al. (2024) and Hu & Li (2024) explored incorporating hierarchical graph information into LLMs. Nevertheless, they mostly focus on the architecture-level incorporation, while we show that it is crucial to integrate the hierarchical information in the instruction tuning data. More importantly, we highlight the consequences of inadequate alignment due to the lack of hierarchical information, i.e., hallucination, and demonstrate the usefulness of the hierarchical information in a wide range of downstream tasks.
3 Graph Tokenization in LGLMs
In this section, we analyze the limitations of node-centric tokenization, which is widely adopted in existing LGLMs.